OpenStudy (anonymous):

what is the most usable half angle formulas in calculus ? i searched the internet for them but they were very complex and never used any of them to solve calculus problems in limit and continuity

OpenStudy (campbell_st):

I always think of half angle formulae as the method of last resort... \[\tan(\frac{\theta}{2}) = t\] this comes from the triangle  and if you use the double angle formulae you can then get \[\tan(\theta) = \frac{2t}{1 -t^2}\] \[\sin (\theta) = \frac{2t}{1 + t^2}\] \[\cos(\theta) = \frac{1 - t^2}{1 + t^2}\] hope this makes some sense...